OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^3 = 216 –6, 3 + 3i *square root* 7, and 3-3i *square root* 7 –6, 3 + 3i *square root* 3, and 3-3i *square root* 3 6, 3 + 3i *square root* 7, and 3-3i *square root* 7 6, –3 + 3i *square root* 3, and -3-3i *square root* 3

OpenStudy (cwrw238):

do you think the real solution is positive or negative?

OpenStudy (anonymous):

positive ?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

so its C or D

OpenStudy (anonymous):

so , it cant be the first two . .

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

c ?

OpenStudy (cwrw238):

x^3 - 216 = (x - 6)(x^2 +12x + 36) = 0 use the quadratic formula to solve the trinomial to get the 2 complex roots

OpenStudy (cwrw238):

sorry the factors are (x - 6)(x^2 +6x + 36) = 0

OpenStudy (cwrw238):

x = [-b +/- sqrt(b^2 - 4ac)] / 2a